OpenStudy (anonymous):

Determine wheter (-9,2) is a solution of 9x+4y=-7

OpenStudy (anonymous):

substitute in the eqyation 9 x -9 + 4 x 2 = -81 + 8 = -73 no

OpenStudy (anonymous):

thank you